OpenStudy (anonymous):

I want to build a right triangular deck behind my house. The hypotenuse of the deck will be 50 feet long. The other two sides of the deck will be equal in length. What is the length of a side of the deck? Round your answer to the nearest hundredth.Answer 35.36 feet 36.78 feet 37.12 feet 38.98 feet

OpenStudy (anonymous):

@bloodrain

OpenStudy (anonymous):

a^2 + b^2 = c^2 a^2 = b^2 c = 50 2a^2 = 2500 a^2 = 1250 a^2 ~ 35.36 So its A) 35.36